OpenStudy (star):

what is the integral of xe^-x with limits 0 to infinity?

myininaya (myininaya):

\[\int\limits_{}^{}xe^{-x}dx=-xe^{-x}-\int\limits_{}^{}-e^{-x}dx=-xe^{-x}-e^{-x}\]

myininaya (myininaya):

before we look at the limits lets check this

myininaya (myininaya):

(-1)e^{-x}--x(-e^{-x})+e^{-x} =xe^{-x} YAY! now for the limists...

myininaya (myininaya):

\[\lim_{b \rightarrow \inf}(-be^{-b}-e^{-b}+e^0)\]

myininaya (myininaya):

e^0=1 -e^{-b}->0 as b->inf now let's look at -be^{-b}

OpenStudy (star):

we are using limits because it is improper integral?

myininaya (myininaya):

yes! :)

OpenStudy (star):

ahh i see. i was wondering where the e^0 came from?

myininaya (myininaya):

i plugged in ther limits the bottom limit is 0 -[0e^0-e^0]=+e^0=+1

myininaya (myininaya):

now we have

myininaya (myininaya):

\[\lim_{b \rightarrow \inf}be^{-b}=\lim_{b \rightarrow \inf} \frac{b}{e^b}=\lim_{b \rightarrow \inf} \frac{1/e^b}\]

myininaya (myininaya):

\[\lim_{b \rightarrow \inf}be^{-b}=\lim_{b \rightarrow \inf} \frac{b}{e^b}\]

myininaya (myininaya):

use l'hospital's rule so we have \[\lim_{b \rightarrow \inf} \frac{1}{e^b}=\lim_{b \rightarrow \inf}e^{-b}=0\]

myininaya (myininaya):

so we have -0-0+1=1

OpenStudy (star):

ah i see! thank you so much for your help :)

myininaya (myininaya):

np